OpenStudy (anonymous):

Construct a subset A of the real numbers such that the set of all limit points of A = the natural numbers

OpenStudy (anonymous):

Construct a subset A of R with \[A' = \mathbb{N} \]

OpenStudy (anonymous):

I THINK I understand the concept of limit points, but getting a function that has limit points which are only natural numbers and not anything else I'm not sure of how to do.

OpenStudy (p0sitr0n):

so you need a bijection from S subset R to N?

OpenStudy (anonymous):

I would think? I never thought of it in terms of bijective. I'm just unsure how to get the set of limit points to contain ONLY natural numbers.

OpenStudy (anonymous):

If your set A were already the natural numbers, then \(A^\prime =\mathbb N \). Since the natural numbers are already closed (with respect to the standard topology on \(\mathbb R\), which I assume is what we are using here).

OpenStudy (anonymous):

Actually, we haven't mentioned anything about closure and topology. Thank you for mentioning that, though, it helps me solve my problem. Is there a way to come to that conclusion without having that knowledge, though?

OpenStudy (anonymous):

@nerdguy2535 Whenever you aren't busy :)

OpenStudy (anonymous):

Hmm, can you tell me the definitions we are working with? That way I can try to help construct a better solution. How are you defining the limit points of a set?

OpenStudy (anonymous):

So, the way we have it defined: "A point p element of the real numbers is said to be a limit point of E if every for every epsilon > 0, the deleted neighborhood of p intersect E is non empty. "A point p element of E is an isolated point of E if there exists epislon > 0 such the epsilon neighborhood p intersect E = {p}" As for other notes we have (sorry if these don't help, thisis just listing all I have to work with) "Let E be a subset of R: a) If p is an element of E', then every neighborhood of p contains infinitely many points of E b) If p is an element of E', then there exists a sequence {p_{n}} in E with p_{n} not equal to p for all n element of the natural numbers, such that p_{n} -> infinity = p" We have that Q' = R. It also saysthat a finite set has no limit points. Then we have an example that proves E' = {0} for E = {1/n | n is an element of N} That's all I have to work with.

OpenStudy (anonymous):

From my understanding, p is a limit point of a set E if there's another point within E for every neighborhood of p. But I guess that hasn't been enough to help me construct a subset per se.

OpenStudy (anonymous):

Alright, this is enough information. Everything is here, you just have to be slightly clever. Do you understand the example that showed E'={0} if \(E=\{\frac{1}{n}\mid n\in \mathbb N\}\)?

OpenStudy (anonymous):

Well....it went through cases where p was 0 (I'm guessing because the example already mentioned 0 was a limit point), where p = 1/n, p was an element of (0,1), p < 0, and p >1. So I suppose all the typical cases you would need. To show p = 0, it just used the archmiedian property, so straight forward. For p = 1/n, it let epsilon = 1/n - 1/(n+1) > 0 and said that the intersection of E and the neighborhood of p = {p}. I'd have to stare at that for a bit, though, lol. For p <0, it said let epsilon = -p > 0 and said that the intersection of the neighborhood of p and E was the empty set. And chose epsilon = p - 1 > 0 for the p > 1 case. For the last 3, it used the neighborhood and not the deleted neighborhood. I guess if you can show something is an isolated point, then that shows it isn't a limit point. But for the last two, saying the intersection of the neighboor of p and E is empty, I'm not sure if that says anythign about the deleted neighborhood without thinking about it more. Either way, this is using a bunch of cases to show 0 is a limit point. I'm not sure how you would do an argument for saying E' is all the natural numbers, though. By contradiction? What about the proof should I be keying in on? xD

OpenStudy (anonymous):

Well, lets not try to jump straight from a set where E'={0} all the way to ALL of the natural numbers. Lets take baby steps. Do you think you could construct a set E where E'={1}? Using the same idea as the one to get E'={0}? (You dont have to write it all out, just tell me if you think you could do it or not).

OpenStudy (anonymous):

Well, I could be wrong, but it looks like you could link finding limit points to finding limits or finding subsequential limits. So if there is that connection, Could I say E = { n/(n+1) | n element of the natural numbers} ? If so, then I would need a case where p = 1, p > 1, and p < 1, right? Would I need a case like in the example where I have something like from (0,1) as well?

OpenStudy (anonymous):

Yep, that would work. I was thinking \(1+\frac{1}{n}\), but your idea works too :) Now how about E'={2}? or E'={3}? How about E'={m} for any natural number m? Do you see the pattern?

OpenStudy (anonymous):

So does that mean n + 1/n works? As a side note, how do you use the equation editor to type 1 + 1/n without making the text go to a separate line? Whenever I click the equation editor, everything I type goes to a different line and makes it all choppy :(

OpenStudy (anonymous):

You don't want \(n+\frac{1}{n}\), since this goes to infinity as n goes to infinity. You want to have another variable m totally separate from n, so something like this: $$m+\frac{1}{n}.$$Then as n goes to infinity, this sequence converges to m. To make equations, type "\ (" "\ )" (without the quotes and space inbetween).

OpenStudy (anonymous):

Thanks ^_^ So if I have something \(m + {1/n}\) that converges to m, what do I have to say about m? Just say that m is an element of n?

OpenStudy (anonymous):

You want to put all these elements together. Create the set: $$E=\{m+\frac{1}{n}\mid m,n\in \mathbb N \}.$$ This set will have all the natural numbers as limit points.

OpenStudy (anonymous):

Alrighty. So, are we assuming m goes to infinity as well, or is it like a constant term and only n goes to infinity?

OpenStudy (anonymous):

While m can go to infinity, that doesn't really help us. You really want to think of it as, "For a fixed m, as n goes to infinity, we have a sequence that converges to m". Since m can be any natural number, we conclude the set of limit points is \(\mathbb N\).

OpenStudy (anonymous):

I see. What concerns me now then is showing they're the only limit points. Like, if m = 1 and n goes to infinity, couldn't you say something like 1.1 was a limit point since you can always find a deleted neighboorhood of 1.1 that contains an element of E for some given n? If that question makes sense.

OpenStudy (anonymous):

Your question does make sense, and it is something we have to be careful about. However, using the exact same ideas from the proof of the E'={0} case, I claim we can indeed show that \(\mathbb N\) are the only limit points.

OpenStudy (anonymous):

So for example, you brought up 1.1. This is \(1+\frac{1}{10}\). This is going to be an isolated point of our set, since the closest points to it are \(1+\frac{1}{9}\) and \(1+\frac{1}{11}\). We will definitely be able to find an epsilon neighborhood around 1.1 that doesn't contain any points from our set in it.

OpenStudy (anonymous):

Pictures really help. Here is the E'={0} case: (Yay for bad pictures). If E is just the set 1/n, it looks something like this. You were able to show that 0 is the only limit point, so all the other points in E are isolted.

OpenStudy (anonymous):

Here is our set: \(E=\{m+\frac {1}{n}\mid m,n\in \mathbb N\}\) It does the same thing as your E' = {0} set, just at every natural number. So its going to behave the exact same way. Every point will be isolated except for the natural numbers.